OpenStudy (anonymous):

I am doing characterstic polynomin and I got -2 (landem) -(landem) square -1.....what would my eigenvectors be?

OpenStudy (anonymous):

sorry its landa..my mistake;)

OpenStudy (anonymous):

see: http://www.sosmath.com/matrix/eigen2/eigen2.html

OpenStudy (anonymous):

solve \[-\lambda^2-2\lambda -1=0\] \[\lambda^2+2\lambda +1=0\] \[(\lambda+1)^2=0\] so one eigenvalue is \[\lambda = -1\]

OpenStudy (anonymous):

then compute \[AX=-X\] or \[AX+X=0\] to get eigenvectors.